OpenStudy (anonymous):

Find the midpoint of the line segment joining the points: (13,6) and (0,6)

OpenStudy (anonymous):

((13+0)/2 , (6+6)/2) (13/2 , 6) is midpoint

OpenStudy (anonymous):

\[(\frac{13}{2},6)\]

OpenStudy (anonymous):

You want to find the midpoint between these points. Therefore find the distance between then in the x and then the y direction. Take half of this and add it to the small points.

OpenStudy (anonymous):

spankss:) do u have a formula for midpoint?

OpenStudy (anonymous):

[((x_1 + x_2)/2) , ((y_1 + y_2)/2)]

OpenStudy (anonymous):

\[(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

ur a lifesavor thanksss!

OpenStudy (anonymous):

u r welcome.